• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 19:26:48 on January23,2002

    Entry number 80425

    keyword=Target cooldown completed

    With help of Vina, we started cool-down after lunch. Loop 2 has a leak and were
    evacuated by Dave and loop 3 is going to be used for LH2.
    Informed MCC at 12:15 to have thme started the procedure.
    At 14:00 started open J-T valve. Most of the tiem J-T at 40-45% flow
    from 10-15 g/s. Diodes are still not working.
    Scattering chamber vacuum was 9.8*E-6 (computer readback was ~9*E-5).
    After cool-down, Vac improved to 2*E-7.
    It was discovered midway that all the CERNOX temperature sensors were not
    working for loop 3. Chris Keith was called in and after checking, it was discovered
    that cables on top of the scattering chamber were broken. Chris and Mike Seely
    are still working to make a new cable to replace the old one. It is estimated to
    take another hour.
    At the mean time, with the help of pressure reading and other temperature indicators
    (VPT and A-B), we completed the cool-down at 19:00 (switched to cold return
    at 18:00, when the inlet temperature went to 25 K and return went up to 130 K!,
    Seems getting worse every time.)
    Since the high power heater is PID on the CERNOX, it is now set at manual.
    When the CERNOX's working again, it should be switch back to PID.